OpenStudy (anonymous):

Jack usually mows his lawn in 6 hours. Marilyn can mow the same yard in 4 hours. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

Let x = the time for both to mow the same lawn working together.\[\frac{1}{6}+\frac{1}{4}\text{=}\frac{1}{x} \]\[\frac{5}{12}\text{= }\frac{1}{x}\]Solve for x. The dimensional unit for x is hours. Another view. \[\frac{1}{6}+\frac{1}{4}=\frac{5}{12} \]Jack and Marilyn can mow 5 lawns in 12 hours. You divide the numerator and denominator by 5 to determine how long it would take to Jack an Marilyn to move the equivalent of one lawn in hours.

OpenStudy (anonymous):

error in the last sentence. The word move should be replaced with the word mow.